The SCC Division of the Big-Drug Company uses standard costing to help control their production costs.  The division’s main product is a pharmaceutical product that comes in a multi-dose package.  The standard variable costs per package are given below:

Direct materials (chemical ingredients 1.2 ounces @$1.00 per ounce)

Direct materials (packaging 1 package piece @$0.40 per package piece)

Direct labor (.20 hour @ $17.00 per hour)

Variable overhead ($2.00 per direct labor hour)

Fixed overhead ($9.00 per direct labor hour)

      During the most recent month, 88,000 finished packages of the product were produced. Other salient information follows:

Actual variable overhead

$35,750.00

Actual fixed overhead

$161,400.00

Actual direct labor hours

17,400

Actual direct labor hourly rate

$17.05

DM chemicals purchase price

$0.980

DM chemicals purchase quantity

115,000

DM packages purchase price

$0.41

DM packages purchase quantity

94,000

DM chemicals used

110,000

DM packages used

92,000

Fixed Overhead Budget

$162,000

For questions 27-30 compute each of the following variances indicating both the dollar amount and whether the amount is favorable (F) or unfavorable (U). If a variance is zero please input zero for the dollar amount and leave the F and U columns blank. DM= direct materials, DL= direct labor.

Dollar Amount

F

U

27.

The DL rate variance

28.

The DL efficiency variance

29.

The variable overhead spending variance

30.

The fixed overhead spending variance

31.  If the fixed overhead is applied based on a rate of $9.00 per direct labor hour, then what is the budgeted fixed overhead base (also called the denominator) quantified in terms of finished packages?
Answer #1
Particulars Variance amount Type Reason
27 DL rate variance = Actual hrs*(Actual rate -Standard rate)
DL rate variance = 17400*(17.05 - 17) $                                   870.00 U It is unfavourable as actual rate paid per hour > standard rate per hour
28 DL effeciency variance = Standard rate*(Standard hrs - Actual hrs)
DL effeciency variance = 17*((88000*0.20)-17400)) $                                3,400.00 F It is favourable variance as actual labour hours < standard labour hours
DL effeciency variance = 17*(17600-17400)
29 Variable OH spending variance = Actual hrs worked*Actual OH rate - Actual hrs worked*standard OH rate
Variable OH spending variance = (35750 - 17400*2) $                                   950.00 U Its is unfavoourable as actual OH > OH that should have been spent at standard rate
30 The fixed overhead spending variance = Budgeted Fixed OH - Actual Fixed OH
The fixed overhead spending variance = 162000 - 161400 $                                   600.00 U It is unfavourable as Actual Fixed OH < Budgeted Fixed OH
31 Calcuulation of Budgeted Fix OH
Fixed overhead ($9.00 per direct labor hour)
Standard hrs required for 1 unit 0.20 hrs
FIXED OH for 1 unit $1.80 ($9*0.20hrs)
Units produced 88000 packages
Budgeted Fixed OH $158,400 ($1.80*88000)